OpenStudy (anonymous):

ok walk me through this please: Find the accumulated amount at the end of 10 years for a principal of $4500 Compounded quarterly at a yearly interest rate of 3%.

OpenStudy (anonymous):

formula is \[4500(1+\frac{.03}{4})^{4\times 10}\]

OpenStudy (anonymous):

which formula is that exactly

OpenStudy (anonymous):

clear what i put where and why?

OpenStudy (anonymous):

ok slowly. principle is 4500 interest is 3%=.03 number of compounding periods per year is 4 number of years is 10

OpenStudy (amistre64):

.03 is the annual rate; but its determined 4 times a year so it gets divided by 4.

OpenStudy (anonymous):

general formula is \[P(1+\frac{r}{n})^{ny}\]

OpenStudy (anonymous):

A=P(1+r)^n

OpenStudy (amistre64):

since the time span is now 4 times ayear; that means that for every year that goes by we have a factor of 4; so ^4t`

OpenStudy (anonymous):

ok is that that formula R=Pi/1-(1+i)^-n

OpenStudy (anonymous):

where r is the interest rate (as a decimal) P is the principle n is the number of compounding periods per year and Y is the number of years

OpenStudy (anonymous):

this is compounded quarterly so you use n = 4

OpenStudy (anonymous):

i get \[4500(1+\frac{.03}{4})^{40}=6067.57\] rounded

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

welcome